# Motility-Driven Viscoelastic Control of Tissue Morphology in Presomitic Mesoderm

### v1

@@ Abstract
During development, embryonic tissues experience mechanical stresses ranging from cellular to supracellular length scales. In response, cells generate active forces that drive rearrangements, allowing the tissue to relax accumulated stresses. The nature of these responses depends strongly on the magnitude and duration of the deformation, giving rise to the tissue’s characteristic viscoelastic behavior. Although experiments have characterized tissue rheology in various contexts, simpler theoretical approaches that directly connect cellular activity to emergent rheological behavior are still limited. In this study, we employ a vertex-based model of epithelial tissue incorporating active force fluctuations in cell vertices to represent cell motility. We capture distinct rounding dynamics and motility-dependent timescales by benchmarking against experimental observations such as the bulging of presomitic mesoderm (PSM) explants driven by Fibroblast Growth Factor(FGF) gradients. Stress relaxation tests reveal rapid short-timescale relaxation alongside persistent long-timescale residual stresses that decrease from anterior to posterior (AP) region of the PSM. By applying oscillatory shear, we analyzed the resulting elastic and viscous responses, revealing motility dependence of storage and loss modulus. Finally, we introduce spatially patterned cues applied in a temporally pulsed manner, mimicking dynamic biochemical or mechanical signals during development. Our results show that while higher motility promotes tissue remodeling in response to these cues, this response is constrained by spatial scale; cellular-scale perturbations are relaxed irrespective of motility strength, preventing complete morphological adaptation.
@@ I Introduction
Embryonic tissues exhibit a complex interplay between mechanical forces and active cellular behaviors during developmentHeisenberg and Bellaïche (2013). In the process, tissue undergoes large-scale deformation and remodeling Mongera et al. (2018). As a response, they propagate forces and relax stresses through active cellular rearrangements and passive mechanical resistance largely governed by cytoskeletal mechanics and actomyosin contractility Kim et al. (2021); Herrera-Perez et al. (2023); Petridou and Heisenberg (2019); Krajnc et al. (2018). The resulting mechanical responses span a wide range of spatial and temporal scales, making tissues highly sensitive to the nature of the applied deformation Fruleux et al. (2024). Understanding these principles not only aids in interpreting experimental measurements of spatial variations in mechanical properties, but also forms the basis for predicting how tissues might reorganize and evolve under different developmental or perturbative conditions.
Tissue behavior on varying space and time scales has been studied extensively for over a decade. Embryonic chicken tissue, such as neural retina, liver, and heart ventricle, has been shown to behave like a viscoelastic material with elastic relaxation on short timescales and a viscous liquid-like response on longer timescales Forgacs et al. (1998). More recently, studies have quantitatively characterized the mechanical properties of embryonic tissues using techniques such as ferrofluid microdroplet deformation Serwane et al. (2017); Mongera et al. (2018), magnetic bead twisting, and micropipette aspiration Campàs et al. (2014). In the presomitic mesoderm (PSM), these approaches have revealed distinct spatial variation in mechanical properties along the anterior–posterior (A–P) axis: both stiffness and viscosity increase from the posterior to the anterior, producing a gradual fluid-to-solid transition across the tissue Mongera et al. (2018); Kim et al. (2021). Importantly, even at a fixed position along this axis, the mechanical response strongly depends on the timescale of deformation Serwane et al. (2017). Localized microdroplet measurements show that tissues initially resist deformation elastically, partially recovering shape on short timescales Mongera et al. (2018). Still, over longer durations, stresses relax through cellular rearrangements and junctional remodeling, leading to a viscous-like flow Mongera et al. (2018); Serwane et al. (2017). Thus, the PSM exhibits a biphasic mechanical response, elastic at short times and viscous at long times. This intrinsic timescale-dependent behavior is a hallmark of tissue viscoelasticity in the developing embryo.
@@
In addition to the macroscopic gradients in mechanical properties observed in the PSM, cellular behaviors also vary systematically along the anterior–posterior (A–P) axis. Time-lapse microscopy shows that cells in the anterior region undergo weaker random spatial motion than those in the posterior region Bénazéraf et al. (2010). This difference in motility mirrors the gradient of fibroblast growth factor (FGF), a key regulator of cell movement Bénazéraf et al. (2010), which decreases from posterior to anterior along the PSM Bénazéraf et al. (2010); Regev et al. (2022). Such coupling between mechanical behavior and cellular activity could be a key determinant of morphogenesis, providing a mechanistic basis for predicting how developing tissues will deform and reorganize under different developmental conditions.
@@
In this work, we use a motile cell-based vertex model Honda and Nagai (2022) to explore how the presomitic mesoderm (PSM) responds to mechanical forces. First, we establish that the model captures key experimental observations, including the characteristic bulging and rounding dynamics of PSM explants in vitro. Using this validated framework, we then extract the viscoelastic timescales that govern tissue mechanics along the anterior–posterior axis. These timescales allow us to make analytical predictions for how the PSM deforms under externally imposed, spatially patterned pulsatile forces. Remarkably, the simulations recapitulate these predictions, providing a direct link between cell motility, viscoelasticity, and large-scale tissue morphology.
@@ II Active Vertex Model
We use the vertex model Honda and Nagai (2022); Fletcher et al. (2014); Farhadifar et al. (2007), a well-established framework for modeling epithelial tissues, to study the emergence of viscoelasticity and its regulation in embryonic PSM. In the vertex model, individual cells are represented as polygons forming a confluent mesh [Figure S1(a)]. The tissue’s mechanical behavior is described in terms of an energy functional that accounts for deviations from target cell geometry and interfacial tensions:
@@
where $\mathbf{r}_{v}$ is the position of vertex $v$, and $\eta$ is an effective friction coefficient that captures viscous resistance from the substrate and/or surrounding fluid. The term $\nabla_{v}U$ represents the gradient of the energy $U$ taken with respect to vertex position. The stochastic term $\boldsymbol{\xi}_{v}(t)$ accounts for active movement of cellular junctions and is modeled as Gaussian white noise with zero mean and amplitude proportional to a motility parameter $\mathcal{M}(\mathbf{r})$ :
@@
where $\delta_{vv^{\prime}}$ ensures that the noise is uncorrelated between different vertices and ${\xi}_{v,\alpha}$ is $x$ or $y$ components of $\boldsymbol{\xi}_{v}$. Together with Eq. 1 and 2, this formulation captures both the mechanical forces that maintain cellular equilibrium and the stochastic, non-equilibrium forces. These stochastic forces effectively represent cellular motility, originating from intracellular processes such as cytoskeletal turnover and actomyosin contractility, and should not be confused with thermal noise Regev et al. (2022). These forces create random movement of cells, which in the context of PSM, parallels Fibroblast Growth Factor (FGF) signaling in coherence with experiments reported in Bénazéraf et al. (2010); Regev et al. (2022) (see Section III for details). Importantly, these random forces at cell vertices drive similar cell–cell contact length fluctuations observed in experiments done on actively fluctuating epithelial junctions Kim et al. (2021).
@@ III Active vertex model reproduces experimental observation involving PSM explants
Figure 1: Effect of motility in presomitic mesoderm (PSM) rheology and relaxation. (a) Change in the shape of the PSM explant over time. Cell motility gradient drives pear-like shape formation by the explant Michaut et al. (2025) (Experimental image is taken with permission from Michaut et al. (2025)). (b) A numerical simulation of the vertex model with a gradient in random cell motility in the form of exponential decay Bénazéraf et al. (2010) from the anterior to posterior axis has been introduced. The model qualitatively reproduces the experimental observation shown in (a). (c) Examples of rounding of anterior and posterior explants from PSM (image taken with permission from Michaut et al. (2025)). Anterior and posterior part shows significant distinction in rounding timescales, suggesting variation in viscosity along the A-P axis. (d) Numerical simulation of the vertex model of rounding of tissue from three different regions, anterior (blue), middle (green), and posterior (red) of the model PSM tissue. A higher cell motility results in faster rounding. (e) Time evolution of circularity of the tissue with three different motility values (as in (d)). (f) Time evolution of overlap function $Q(t)$ (see Equation S7). It shows the rate of radial movement of a cell from its initial position. The dotted line marks when $Q(t)$ drops below $1/e$ of its initial value. (g) Dependence of viscosity $\eta$, calculated from the Green-Kubo relation (see Materials and Methods for definition), and $\alpha$ relaxation Time $\tau_{\alpha}$ on motility. The anterior region shows a high value of viscous timescales, and it drops linearly towards the posterior direction.
To benchmark our simulation parameters ($\lambda$, $\beta$, $\gamma$) with PSM, we have reproduced an experimental result reported in Michaut et al. (2025), in which an embryonic PSM cultured ex vivo exhibits spontaneous shape remodeling. Experimental studies have shown that isolated PSM explants spontaneously remodel their shape and fluidize over time, suggesting a tight coupling between motility and mechanical relaxation. By tuning our parameters to capture these behaviors, we establish a consistent framework for exploring how spatial variations in motility and mechanical resistance along the anterior–posterior axis shape tissue morphogenesis.
@@ III.1 Cell motility drives bulging of PSM explant
The PSM exhibits a gradient in cellular motility along its anterior-posterior axis, regulated by Fibroblast Growth Factor (FGF) signaling Bénazéraf et al. (2010); Regev et al. (2022). The highly motile cells in the posterior PSM are able to overcome intrinsic contractile and adhesive forces and adapt a circular shape [Fig. 1(a)]. In contrast, the anterior region, with lower motility, fails to circularize, resulting in a characteristic pear-like tissue shape [Fig. 1(a)]. Notably, this morphological transformation occurs on timescales shorter than those of cell division Michaut et al. (2025). To gain deeper insight into the spatial variation of mechanical properties along the anterior-posterior (AP) axis, Arthur et al.Michaut et al. (2025) isolated tissues from different regions of the presomitic mesoderm (PSM) and allowed them to evolve freely. While both explants ultimately adopted a circular shape, they did so over significantly different timescales [Fig.1(c)]. This rounding up of a tissue is a hallmark of fluidization as fluids try to minimize the surface energy by forming a circular shape. These experiments reveal that motility can drive tissue fluidization, which can control PSM morphogenesis.
We have reproduced these experiments using the vertex model where starting from a rectangular tissue, we incorporate FGF gradient by making the strength of motility $\left(\mathcal{M}(\textbf{r})\right)$ decay exponentially from posterior to anterior direction with a length scale established by experimental observations Bénazéraf et al. (2010); Regev et al. (2022); Michaut et al. (2025). To be consistent with the experiments, we start with an initial tissue state that, due to contractility and the adhesive nature of the cells, shrinks in size initially, but over longer times, motile cells easily overcome the energy barrier required for junctional rearrangement, which facilitates fluidization. This motility-driven fluidization varies along the A-P axis. The posterior PSM being in a higher state of motility becomes rounded in shape quicker than the anterior part, ultimately leading to the emergence of the characteristic pear-like morphology with a decrease in total tissue length[Fig. 1(b)]. We also simulated tissue evolution independently under three different motility strengths. Consistent with experimental findings, we observe that tissues mimicking the anterior region reach circularity more rapidly than those representing the middle or posterior PSM [Fig. 1(d,e)].
@@
In our simulations, spreading arises from motility-driven junctional rearrangements, especially T1 transitions, where cells exchange neighbors. These rearrangements allow local structural reorganization. In regions with higher motility, such as the posterior PSM, cells more frequently overcome the energetic barriers associated with junctional remodeling, resulting in a more fluid-like behavior (Fig. S3). This inhomogeneous enhancement in tissue fluidity promotes inhomogeneous tissue deformation that causes the reproduction of experimentally observed morphology. Our results thus demonstrate active fluctuation driven rearrangements are sufficient to drive large-scale tissue remodeling in the absence of proliferation.
@@ III.2 Motility gradient leads to rheological heterogeneity along PSM
Experimental studies have suggested that variations in tissue viscosity along the A-P axis play a key role in shaping the morphogenesis of the PSM Mongera et al. (2018); Michaut et al. (2025). To measure this rheological heterogeneity in our simulations, we quantified spatial variations in emergent tissue viscosity using two different approaches: the Green-Kubo formalism (see S.I., Sec. S2.3.2) and the overlap function approach (S.I. Sec. S2.2.1). The Green-Kubo formalism shows that the anterior region of the PSM exhibits significantly higher viscosity compared to the posterior, with viscosity decreasing along the AP axis (Fig. 1(g)). These results are in strong qualitative agreement with experimental measurements reported in Ref.Michaut et al. (2025). Further analysis of cell dynamics across three distinct regions reveals that the overlap function $Q(t)$, which measures the radial displacement of cells from their initial positions, exhibits three quantitatively distinct temporal decay profiles (Figure 1(f)). From the decay of $Q(t)$, we estimate the relaxation time $\tau_{\alpha}$ [see Materials and Methods for definition], which provides an estimate of the timescale over which tissue rearranges. We find that $\tau_{\alpha}$ is shortest in the posterior region and longest in the anterior region [Fig. 1(g)], confirming that motility-driven fluidization is most prominent toward the posterior PSM. This gradient in viscous relaxation timescales reinforces the conclusion that tissue fluidity is spatially regulated and correlates with the observed morphological dynamics.

### v2

@@ Abstract
Embryonic tissues deform across broad spatial and temporal scales and relax stress through active rearrangements. A quantitative link between cell-scale activity, spatial forcing, and emergent tissue-scale mechanics remains incomplete. Here, we use a vertex-based tissue model with active force fluctuations to study how motility controls viscoelastic response. After validation against experimental presomitic mesoderm relaxation dynamics, we extract intrinsic mechanical timescales using stress relaxation and oscillatory shear. The model captures motility-dependent shifts between elastic and viscous behavior and the coexistence of fast relaxation with long-lived residual stress. When subjected to spatially patterned, temporally pulsed forcing, tissues behave as mechanical filters: long-wavelength inputs are accumulated, whereas short-wavelength, cell-scale perturbations are rapidly erased, largely independent of motility. Simulations with localized motility hotspots, motivated by spatially confined FGF signaling reported in vertebrate limb development, produce sustained protrusive tissue deformations consistent with experimentally observed early bud-like morphologies. Together, these results establish a minimal framework linking motility-driven activity to wavelength-selective mechanical memory and emergent tissue patterning.
@@ I Introduction
Embryonic tissues operate far from mechanical equilibrium, where active cellular processes and passive material responses are tightly intertwined heisenberg2013forces. During development, tissues undergo large-scale deformation and remodeling mongeraFluidtosolidJammingTransition2018. This remodeling occurs via active cellular rearrangements and passive mechanical resistance, largely governed by cytoskeletal mechanics and actomyosin contractility kimEmbryonicTissuesActive2021a; PRXLife.1.013004; petridou2019tissue; krajncFluidizationEpithelialSheets2018. These mechanical responses span a wide range of spatial and temporal scales, making tissues highly sensitive to the nature of the applied deformation fruleuxGrowthCouplesTemporal2024. Understanding the physical principles of this mechanoregulation is essential not only for interpreting spatially heterogeneous mechanical measurements but also for predicting how tissues reorganize and adapt under developmental and perturbative conditions.
Towards this, advanced experimental approaches, such as ferrofluid microdroplet deformation serwaneVivoQuantificationSpatially2017a; mongeraFluidtosolidJammingTransition2018, magnetic bead twisting, and micropipette aspiration campas2014quantifying, have enabled direct estimation of tissue rheology. Using these works, the viscoelastic nature of different embryonic tissues has been established, where elastic and viscous responses are dominant on shorter and longer timescales, respectively forgacsViscoelasticPropertiesLiving1998. For example, in the presomitic mesoderm (PSM) in vertebrates, a spatial gradient of mechanical properties exists along the anterior–posterior (A–P) axis, with both tissue stiffness and viscosity increasing from the posterior to the anterior end. This static pattern results in a posterior-to-anterior gradient from a more fluid-like to a more solid-like state across the tissue mongeraFluidtosolidJammingTransition2018; kimEmbryonicTissuesActive2021a. Importantly, even at a fixed position along this axis, the mechanical response depends strongly on the timescale of deformation- elastic at short time scales and viscous at large timescales mongeraFluidtosolidJammingTransition2018; serwaneVivoQuantificationSpatially2017a. This intrinsic timescale-dependent behavior is a hallmark of tissue viscoelasticity in the developing embryo.
@@
These tissue-scale viscoelasticity gradients in vertebrate PSM are accompanied by variations in spatial motility of the cells. Time-lapse microscopy shows that cells in the anterior region undergo weaker random spatial motion than those in the posterior region benazerafRandomCellMotility2010. This motility gradient closely correlates with the spatial distribution of fibroblast growth factor (FGF), a central regulator of cell movement, whose expression level decreases from posterior to anterior benazerafRandomCellMotility2010; regev2022rectified. The coexistence of spatially varying motility and mechanical properties suggests a tight coupling between active cellular processes and emergent tissue rheology, providing a potential mechanistic basis for how developing tissues deform, reorganize, and pattern under physiological conditions.
Another aspect of tissues with graded mechanical properties is how cells perceive and respond to dynamic forces. During embryogenesis, for example, tissues experience intrinsic and extrinsic pulsatile cues across wide spatiotemporal scales. Intrinsically, key morphogenetic events are driven by periodic cellular activities: actomyosin pulses produce ratcheted apical constrictions sutherland2020pulsed, pulsed contractions power dorsal closuresolon2009pulsed, and oscillatory gene expression patterns direct boundary formation during somitogenesis hubaud2014signalling. Extrinsically, rhythmic forces from neighboring organs also exert widespread influence. Early heartbeats, initiated before functional circulation, generate pulsatile pressure waves that propagate through millimeter-scale tissues and can mechanically stimulate adjacent structures such as the presomitic mesoderm hornberger2007rhythm; watanabe2014regular; tyser2020first. Meanwhile, cilia-driven left–right nodal flow creates oscillatory shear stress that not only breaks bilateral symmetry but may also deliver mechanical stimuli to nearby mesodermal regions shinohara2017cilia. Collectively, these cyclic mechanical inputs help locally modulate cell behavior and globally orchestrate tissue-level morphogenetic processes. Thus, mechanical heterogeneity within developing tissues may serve as a critical modulator of rhythmic force transmission, enabling the spatiotemporal coordination of morphogenesis.
@@
Despite substantial experimental progress, a predictive framework that connects active cellular motility, tissue viscoelasticity, and the spatiotemporal structure of mechanical cues remains lacking. In embryonic tissues, where pronounced gradients exist in both mechanical properties and cellular activity, it remains unclear how these features jointly determine whether a dynamic mechanical input is filtered, propagated, or permanently encoded as a tissue-scale deformation. Here, we bridge this gap using a motile cell–based vertex model honda2022vertex to investigate the mechanical response of the presomitic mesoderm. We first establish that the model reproduces key experimental signatures, including the characteristic bulging and rounding dynamics of PSM explants in vitro. Leveraging this validated framework, we extract the viscoelastic timescales that govern tissue mechanics along the anterior–posterior axis and develop analytical predictions for deformation under spatially patterned, pulsatile forcing. We also perform simulations to recapitulate these predictions, directly linking cell-scale motility, emergent viscoelasticity, and large-scale morphogenetic shape change.
@@ II Active Vertex Model
We use the vertex model honda2022vertex; fletcher2014vertex; farhadifar2007influence, a well-established framework for modeling epithelial tissues, to study the emergence of viscoelasticity and its regulation in embryonic PSM. In the vertex model, individual cells are represented as polygons forming a confluent mesh [Figure S1(a)]. The tissue’s mechanical behavior is described in terms of an energy functional that accounts for deviations from target cell geometry and interfacial tensions:
@@
where $\mathbf{r}_{v}$ is the position of vertex $v$ [Figure S1(a)], and $\eta$ is an effective friction coefficient that captures viscous resistance from the substrate and/or surrounding fluid. The term $\nabla_{v}U$ represents the gradient of the energy $U$ taken with respect to vertex position. The stochastic term $\boldsymbol{\xi}_{v}(t)$ accounts for active movement of cellular junctions and is modeled as Gaussian white noise with zero mean and amplitude proportional to a motility parameter $\mathcal{M}(\mathbf{r})$ :
@@
where $\delta_{vv^{\prime}}$ ensures that the noise is uncorrelated between different vertices and ${\xi}_{v,\alpha}$ is $x$ or $y$ components of $\boldsymbol{\xi}_{v}$. Together with Eq. 1 and 2, this formulation captures both the mechanical forces that maintain cellular equilibrium and the stochastic, non-equilibrium forces. These stochastic forces effectively represent cellular motility, originating from intracellular processes such as cytoskeletal turnover and actomyosin contractility, and should not be confused with thermal noise regev2022rectified. These forces create random movement of cells, which in the context of PSM, parallels fibroblast growth factor (FGF) signaling in coherence with experiments reported in benazerafRandomCellMotility2010; regev2022rectified (see Section III for details). Importantly, these random forces at cell vertices drive similar cell–cell contact length fluctuations observed in experiments done on actively fluctuating epithelial junctions kimEmbryonicTissuesActive2021a.
@@ III Active vertex model reproduces experimental observation involving PSM explants
Figure 1: Effect of motility in presomitic mesoderm (PSM) rheology and relaxation. (a) Change in the shape of the PSM explant over time. Cell motility gradient drives pear-like shape formation of the explant (Experimental image is taken with permission from MICHAUT2025). (b) A numerical simulation of the vertex model with a gradient in random cell motility in the form of exponential decay benazerafRandomCellMotility2010 from the posterior to anterior tissue has been introduced. The model qualitatively reproduces the experimental observation shown in (a). (c) Examples of rounding of anterior and posterior explants from PSM (image taken with permission from MICHAUT2025). The Anterior and posterior parts show significant distinction in rounding timescales, suggesting variation in viscosity along the A-P axis. (d) Numerical simulation of the vertex model of rounding of tissue from three different regions, anterior (blue), middle (green), and posterior (red) of the model PSM tissue. A higher cell motility results in faster rounding. (e) Time evolution of circularity of the tissue with three different motility values (as in (d)). (f) Time evolution of overlap function $Q(t)$ (see Equation S7). It shows the rate of radial movement of a cell from its initial position. The dotted line marks when $Q(t)$ drops below $1/e$ of its initial value. (g) Dependence of viscosity $\eta$, calculated from the Green-Kubo relation (see S.I. Sec. S2.3.2 for definition), and $\alpha$ relaxation Time $\tau_{\alpha}$ on motility. The anterior region shows a high value of viscous timescales, which decreases towards the posterior direction.
To benchmark our simulation parameters ($\lambda$, $\beta$, $\gamma$) against PSM behavior, we reproduce an experimental result from MICHAUT2025, where an embryonic PSM cultured ex vivo undergoes spontaneous shape remodeling. Experiments show that isolated PSM explants remodel their shape and fluidize over time MICHAUT2025, demonstrating a tight coupling between motility and mechanical relaxation. By tuning our parameters to capture these behaviors, we establish a consistent framework for analyzing how spatial variations in motility and mechanical resistance along the anterior–posterior axis govern tissue morphogenesis.
@@ III.1 Cell motility drives bulging of PSM explant
The PSM exhibits a graded profile of cellular motility along its anterior–posterior (AP) axis, regulated by fibroblast growth factor (FGF) signaling benazerafRandomCellMotility2010; regev2022rectified. Highly motile cells in the posterior PSM can overcome intrinsic contractile and adhesive forces, enabling the explant to adopt a circular shape [Fig. 1(a)]. By contrast, the less motile anterior region fails to circularize, producing the characteristic pear-like morphology of the tissue [Fig. 1(a)]. This shape transformation occurs on timescales shorter than those of cell division MICHAUT2025, indicating that it is driven by mechanics rather than growth.
To probe how mechanical properties vary along the AP axis, Arthur et al. MICHAUT2025 isolated tissues from different PSM regions and monitored their free evolution. Both anterior and posterior explants ultimately rounded into circular shapes, but on markedly different timescales [Fig. 1(c)]. Such rounding is a hallmark of fluidization, as a fluid minimizes surface energy by approaching a circular geometry. These observations demonstrate that motility-induced fluidization governs PSM morphogenesis.
We reproduce these experimental observations using an active vertex model. Starting from a rectangular tissue, we implement the FGF gradient by prescribing an exponentially decaying motility field, $\mathcal{M}(\mathbf{r})$, decreasing from posterior to anterior with a characteristic length scale set by experimentally measured motility decay benazerafRandomCellMotility2010; regev2022rectified; MICHAUT2025. To remain consistent with experiments, we initialize the tissue in a state that initially contracts due to cellular contractility and adhesion. At longer times, motile cells overcome the energy barriers associated with junctional rearrangements, enabling fluidization. Because motility varies along the AP axis, fluidization occurs heterogeneously: the posterior PSM, with higher motility, rounds rapidly, whereas the anterior region remains elongated longer, ultimately producing the characteristic pear-like morphology accompanied by a net decrease in tissue length [Fig. 1(b)].
We further simulated tissue evolution under three distinct motility strengths. In agreement with experimental trends, tissues corresponding to posterior PSM reach circularity significantly faster than those representing middle or anterior regions [Fig. 1(d,e)].
@@
In our simulations, tissue spreading emerges from motility-driven junctional rearrangements, particularly T1 transitions in which cells exchange neighbors. These rearrangements enable local structural reorganization. In regions with higher motility, such as the posterior PSM, cells more readily overcome the energetic barriers associated with junctional remodeling, leading to more fluid-like behavior (Fig. S3). This spatially heterogeneous increase in tissue fluidity drives nonuniform deformation, reproducing the experimentally observed morphology. Our results thus demonstrate that active, fluctuation-driven rearrangements are sufficient to induce large-scale tissue remodeling even in the absence of cell proliferation.
@@ III.2 Motility gradient leads to rheological heterogeneity along PSM
Experimental studies suggest that variations in tissue viscosity along the AP axis play a key role in shaping PSM morphogenesis mongeraFluidtosolidJammingTransition2018; MICHAUT2025. To quantify this rheological heterogeneity in our simulations, we measured spatial variations in emergent tissue viscosity using two complementary approaches: the Green–Kubo formalism (S.I., Sec. S2.3.2) and the overlap function method (S.I., Sec. S2.2.1). The Green–Kubo analysis shows that the anterior PSM exhibits substantially higher viscosity than the posterior, with viscosity decreasing smoothly along the AP axis [Fig. 1(g)]. These results are in strong qualitative agreement with experimental measurements reported in  MICHAUT2025. To further probe dynamical heterogeneity, we examined cell motions across three representative regions using the overlap function $Q(t)$, which quantifies the radial displacement of cells from their initial positions. The resulting decay profiles display clear quantitative differences among anterior, middle, and posterior regions [Fig. 1(f)]. From these curves, we extracted the relaxation time $\tau_{\alpha}$ (see S.I. Sec. S2.2.1), which characterizes the timescale for structural rearrangements. We find that $\tau_{\alpha}$ is shortest in the posterior and longest in the anterior PSM [Fig. 1(g)], indicating that motility-driven fluidization is most pronounced toward the posterior. This gradient in viscous relaxation timescales confirms that tissue fluidity is spatially regulated and correlates with the observed morphological dynamics.